OpenStudy (anonymous):

Aimee and Ben are purchasing a condominium and are financing $415,500. The mortgage is a 20-year 3/1 ARM at 2.75% with a cap structure of 1.5/8. What will their payments be for the first 3 years? $1,696.24 $2,252.70 $2,572.92 $3,475.41

OpenStudy (anonymous):

do any of you know this topic?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

prop 3

OpenStudy (anonymous):

how do you know @beomsu.c.59131

OpenStudy (tkhunny):

Well, for a 20 year 3/1 ARM, you can just ignore that it's an ARM for the first 3 year. Calculate the payment for a 20 year fixed 2.75%.

OpenStudy (tkhunny):

Why not ask the right answer how instead of the two wrong answers?

OpenStudy (anonymous):

so how is c the answer @tkhunny

OpenStudy (tkhunny):

It isn't. Why do you think it is? Did you calculate it, ignore that it's an ARM?

OpenStudy (anonymous):

but what equation do i use @tkhunny it would the monthly payment right?

OpenStudy (tkhunny):

Yes. Let's see it.

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Not sure what that formula is. i = 0.0275 m = 12 j = i/m = 0.0022916666666666... v = 1/(1+j) = 0.997713573061734... n = 20 (years) \(Pmt = \dfrac{415500}{\dfrac{v\left(1 - v^{n*m}\right)}{1 - v}}\) That can be algebraically simplified a little, but that's it.

OpenStudy (anonymous):

ok but i put the formula there @tkhunny

OpenStudy (tkhunny):

I still don't know what formula you posted. Define the terms and maybe we can make some sense of it.

OpenStudy (anonymous):

It is the monthly payment formula

OpenStudy (anonymous):

I solved it with your equation and I D? IS THAT RIGHT @tkhunny

OpenStudy (tkhunny):

You are going to have to show your work. I can't tell what you are doing if you don't show me. You still didn't define the terms in your monthly payment formula. Did you see how cleanly I laid them out?

OpenStudy (anonymous):

Ok but do you want the definition but did i get it right at least?

OpenStudy (tkhunny):

You need to tell me if it is right. Confidence is part of what we are doing, here.

OpenStudy (anonymous):

Yes I do feel it is right! @tkhunny I am just asking you because you have more knowledge than I do on the subject.

OpenStudy (tkhunny):

Well, I do like confidence, but you also have to be right. Show your work. Let's see where you wandered off. Notice how I defined v in terms of the MONTHLY interest.

OpenStudy (anonymous):

well i just plugged this in the calculator @tkhunny

OpenStudy (tkhunny):

There is no such thing as "just plugged". You did it in a certain order and you made certain assumptions. Just show your work. Did you get the same j and v that I did? If you aren't using my formula, then show the results of your formula after defining the terms for me.

OpenStudy (anonymous):

@Jesstho.-.

OpenStudy (anonymous):

B